OpenStudy (anonymous):

Divide and simplify. h/b^2 ÷ h^2/b^3

OpenStudy (anonymous):

factor and cancel.

OpenStudy (saifoo.khan):

\[\frac{h}{b^2} \times \frac{b^3}{h^2}\]

OpenStudy (anonymous):

h/b^2 div h^2/b^3 h/b^2 * b^3/h^2 b/h

OpenStudy (anonymous):

or in this case just cancel

OpenStudy (saifoo.khan):

\[\frac{b^{3-2}}{h^{2-1}}\]

OpenStudy (saifoo.khan):

\[\frac{b}{h}\]

OpenStudy (anonymous):

\[\frac{\displaystyle\frac{h}{b^2}}{\displaystyle\frac{h^2}{b^3}} = (h\cdot b^{-2})\cdot(h^2\cdot b^{-3})^{-1} = (h \cdot b^{-2})\cdot (h^{-2}b^3) = (h\cdot h^{-2})\cdot(b^3\cdot b^{-2}) = h^{-1}\cdot b = \frac{b}{h} \]